A solution was prepared by dissolving 39.0g of KCl in 225 g of water.

1. Calculate the mass percent of KCl in the solution.

2. Calculate the mole fraction of KCl in the solution.

3. Calculate the molarity of KCl in the solution if the total volume of the solution is 239 mL.

4. Calculate the molality of KCl in the solution.
Answer #1

39+22) )234 12-5

Moluv mc Soin in utres O J23 2、188 M de 0.623 O22J
Answer #2

. % = (39.0 / 39.0 + 225 ) x 100 = 14.8 %

2.
Molecular weight KCl = 74.55 g/mol
39.0 / 74.55 = 0.523 moles KCl
Molecular weight H2O = 18 g/mol
225 / 18 = 12.5 moles H2O
X (KCl) = 0.523 / 0.523 + 12.5 = 0.0401

3.
V = 0.239 L
M = 0.523 / 0.239 = 2.19

4.
m = 0.523 /0.225 Kg = 2.32
Answer #3

1. % = (39.0 / 39.0 + 225 ) x 100 = 14.8 %

2.
Molecular weight KCl = 74.55 g/mol
39.0 / 74.55 = 0.523 moles KCl
Molecular weight H2O = 18 g/mol
225 / 18 = 12.5 moles H2O
X (KCl) = 0.523 / 0.523 + 12.5 = 0.0401

3.
V = 0.239 L
M = 0.523 / 0.239 = 2.19

4.
m = 0.523 /0.225 Kg = 2.32
Answer #4

A solution was prepared by dissolving 39.0g of KCl in 225 g of water.

1. Calculate the mass percent of KCl in the solution.